OpenStudy (anonymous):

Hey I have a problem with limits. In my example, I have: lim(sinx * cosx - sinx)/x^3 , as x->0 then it becomes equal to: lim(sinx/x) * (cosx-1)/x^2 My question is: how did he go from sinx * cosx - sinx to cosx - 1. I'm thinking it's an identity but cannot find it. Any insight would be much appreciated.

OpenStudy (anonymous):

\[\lim \frac{ \sin(x) \cos(x) - \sin(x) }{ x^{3} } \] as x-> 0

OpenStudy (anonymous):

to: \[\lim \frac{ \sin(x) }{ x } * \frac{ \cos(x)-1 }{ x^{2} }\]

OpenStudy (anonymous):

How did my example jump to this equation?

OpenStudy (psymon):

All that was done is a sin was factored out. sinxcosx - sinx just became sinx(cosx-1). As for the splitting of fractions, theres no identity, its just a visual. It was written ina different way so you can maybe see whats going on. Because for example, sinx/x as x goes to 0 is an identity and it equals 1. But you might not see that unless its written in that way.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

OMG thanks so much psymon hahahahaha i love you!

OpenStudy (psymon):

yeah, np :3